### v1

@@ 3.6 Heterogeneous Layer Assignment Algorithm
[…]
9:if $M_{\text{embed}}\leq M_{\text{embed}}^{\max}$ then
10:  $\mathcal{A}\leftarrow\{(\text{Embed},D_{\text{embed}}),(\text{LMHead},D_{\text{embed}})\}$
11:  Compute: $E_{\text{embed}}\leftarrow P_{\text{embed}}\times(L\times d_{h}/B_{\text{embed}})$
12:else
13:  Raise InsufficientMemoryError
14:end if
15:Step 3: Greedy Decoder Assignment
[…]
17:for each decoder layer $l\in[1,n_{d}]$ do
18:  $D_{\text{current}}\leftarrow\mathcal{D}_{\text{sorted}}[\text{device\_idx}]$
19:  $M_{l}\leftarrow(2d_{h}^{2}+8d_{h}^{2}+2Td_{h})\times 2$ bytes
20:  while $M_{l}>M_{\text{current}}^{\max}$ available do
21:   $\text{device\_idx}\leftarrow\text{device\_idx}+1$
22:  end while
23:  Compute: $E_{\text{prefill},l}\leftarrow P_{\text{current}}\times\frac{4Td_{h}^{2}}{C_{\text{current}}}$
24:  Compute: $E_{\text{decode},l}\leftarrow P_{\text{current}}\times(S-1)\times\frac{2Td_{h}^{2}}{B_{\text{current}}}$
25:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{(\text{Layer}_{l},D_{\text{current}})\}$; $E_{\text{total}}\leftarrow E_{\text{total}}+E_{\text{prefill},l}+E_{\text{decode},l}$
26:end for
27:Step 4: Constraint Validation
28:if $\tau_{\text{total}}>\tau_{\max}$ or $E_{\text{total}}>E_{\max}$ then
29:  Adjust $S$ or apply quantization; retry
30:end if
31:return $\mathcal{A}$, $E_{\text{total}}$, $\tau_{\text{total}}$
3.6.1 Energy Calculation Helper Functions
For each layer-device assignment, energy consumption is calculated by decomposing the inference into prefill and decode phases, each exhibiting different computational and memory access characteristics.
Prefill Phase Energy (Prompt Processing): During prefill, all input tokens are processed simultaneously, exposing high arithmetic intensity:
$E_{\text{prefill}}=P_{i}\cdot\tau_{\text{prefill}}=P_{i}\cdot\frac{\text{FLOPs}_{\text{prefill}}}{C_{i}}$ (19)
where the FLOP count per layer for prefill is:
$\text{FLOPs}_{\text{prefill}}=4\times T\times d_{h}^{2}$ (20)
This accounts for: (1) multi-head attention computing $\mathbf{Q}\mathbf{K}^{\top}$ and value aggregation (cost $2T^{2}d_{h}$, but amortized to $2Td_{h}^{2}$ with KV cache projection), and (2) feed-forward network with intermediate dimension expansion ($2\times d_{h}\times 4d_{h}=8d_{h}^{2}$ FLOPs). The latency is bound by compute capacity $C_{i}=2f_{i}\cdot n_{\text{cores}}$ (FLOPs per second).
Decode Phase Energy (Autoregressive Generation): Each subsequent token is processed autoregressively (one-at-a-time), exposing memory-bound characteristics:
$E_{\text{decode}}=P_{i}\cdot\tau_{\text{decode}}=P_{i}\cdot(S-1)\cdot\frac{\text{Bytes}_{\text{decode}}}{B_{i}}$ (21)
where the memory bandwidth constraint dominates (not compute):
$\text{Bytes}_{\text{decode}}=2\times T\times d_{h}+2\times d_{h}^{2}$ (22)
This accounts for: (1) loading KV cache ($2\times T\times d_{h}$ bytes), and (2) loading layer weights for computation ($2\times d_{h}^{2}$ bytes for projection matrices). The latency is bound by memory bandwidth $B_{i}$ (bytes per second), since the arithmetic intensity is low.
3.6.2 Efficiency Factor and Power Estimation
The Intelligence Per Watt (IPW) metric for a given allocation is computed as:
$\text{IPW}_{\text{allocation}}=\frac{C(S,N,T)}{P_{\text{avg}}}$ (23)
where $C(S,N,T)$ is coverage (Theorem 1.1, Eq. 1) and the average power during inference is:
$P_{\text{avg}}=\frac{1}{\tau_{\text{total}}}\sum_{i}P_{i}\cdot\tau_{i}$ (24)
This is the time-weighted average of device power draws. For example, if prefill takes 100ms on a GPU (300W) and decode takes 1000ms on an NPU (10W), the average power is $(300\times 0.1+10\times 1.0)/1.1=38.2$W.
3.6.3 Latency Estimation and Sequencing
Total inference latency is decomposed into four components reflecting different bottlenecks:
$\tau_{\text{total}}=\tau_{\text{prefill}}+\tau_{\text{decode}}+\tau_{\text{io}}+\tau_{\text{overhead}}$ (25)
Prefill Latency (Bottleneck Across Devices):
$\tau_{\text{prefill}}=\max_{i}\sum_{l\in L_{i}}\frac{4T\cdot d_{h}^{2}}{C_{i}}$ (26)
where the max is taken over all devices (the slowest device determines total prefill time due to sequential dependencies). If prefill is split across multiple devices, the critical path (slowest device) dominates.
Decode Latency (Sequential Per-Sample):
$\tau_{\text{decode}}=(S-1)\cdot\max_{i}\left(\sum_{l\in L_{i}}\frac{2\times T\times d_{h}^{2}}{B_{i}}+\text{overhead}_{i}\right)$ (27)
where decode proceeds sequentially token-by-token (each of S1S-1 subsequent tokens requires a full forward pass through all layers).
I/O and Data Movement Overhead:
$\tau_{\text{io}}=\sum_{(i,j):\text{adjacent devices}}\frac{T\times d_{h}\times 2\text{ bytes}}{\text{BW}_{\text{interconnect}}}$ (28)
When consecutive layers are assigned to different devices (e.g., embedding on CPU, decoder on GPU), intermediate activations must be transferred. The $T\times d_{h}$ values represent the hidden state dimension; the factor of 2 accounts for FP16 (2 bytes per value).
Scheduling Overhead:
$\tau_{\text{overhead}}=\text{const}+\alpha\cdot\log(S)$ (29)
In heterogeneous systems, task scheduling, kernel launch delays, and memory allocation add latency. The constant term (typically 1-5ms) reflects fixed setup costs; the logarithmic term reflects queue depth overhead that grows sublinearly with sample count.
3.7 Dynamic Runtime Orchestration and Adaptive Execution
At inference time, the framework executes the pre-computed allocation plan with adaptive sample budget adjustment based on real-time energy and latency monitoring:
Algorithm 2 QEIL: Runtime Orchestration
0: Allocation plan $\mathcal{A}$; Input tokens $\mathbf{x}$; Energy budget $E_{\max}$
0: Inference output with adaptive samples
1:Phase 1: Prefill
2:for each device $D_{i}$ with prefill layers do
3:  Execute embedding and decoder prefill on $D_{i}$
4:  Record $\tau_{\text{prefill},i}$, $E_{\text{prefill},i}$
5:end for
6:Phase 2: Adaptive Decode
[…]
8:while $\text{sample\_idx}<S$ AND $(E_{\text{current}}+E_{\text{decode\_est}})<E_{\max}$ do
9:  for each device $D_{i}$ with decoder layers (sequential) do
10:   Execute decoder on $D_{i}$ for current token
11:   $E_{\text{current}}\leftarrow E_{\text{current}}+E_{\text{decode},i}$
12:  end for
13:  Project to logits via LM Head; sample token
14:  $\text{sample\_idx}\leftarrow\text{sample\_idx}+1$
15:end while
16:return Output tokens, $E_{\text{current}}$, $\tau_{\text{actual}}$
3.8 Inference Efficiency Metrics
QEIL employs three complementary metrics that capture different aspects of edge deployment, each optimizing for different hardware and power constraints:
3.8.1 Intelligence Per Watt (IPW)
Measures steady-state power efficiency, capturing how much task accuracy is achieved per unit instantaneous power:
$\text{IPW}=\frac{C(S,N,T)}{P_{\text{avg}}}\quad[\text{coverage/watt}]$ (30)
where $C(S,N,T)$ is coverage (Theorem 1.1, Eq. 1) and $P_{\text{avg}}$ is average power consumed during inference. IPW directly captures task accuracy per unit power, making it most relevant for battery-operated devices where power draw determines battery lifetime. A device with IPW = 0.5 coverage/watt achieves 50% coverage with 1 watt power draw.
3.8.2 Energy-Coverage Efficiency (ECE)
Measures total energy consumed per unit coverage, accounting for both power draw and latency:
$\text{ECE}=\frac{C(S,N,T)}{E_{\text{total}}/1000}\quad[\text{coverage/joule}]$ (31)
where $E_{\text{total}}$ is the cumulative energy in joules. ECE optimizes for energy budget constraints (e.g., ”solve as many queries as possible with 100kJ”) rather than power constraints. A device with ECE = 2 coverage/joule achieves full coverage (100% accuracy) using 50 joules of energy.
3.8.3 Price-Power-Performance (PPP) Score
Dimensionless composite metric balancing economic cost, power consumption, and throughput:
$\text{PPP}=\frac{\text{Throughput}_{\text{queries/sec}}\times C(S,N,T)}{P_{\text{avg}}\times\text{Cost}_{\text{amortized/query}}}$ (32)
PPP captures the total return-on-investment: queries solved per second per watt per dollar. This metric enables fair comparison across heterogeneous configurations with different hardware prices (cheap old GPUs vs. expensive new NPUs), power characteristics (CPUs consuming 45W vs. NPUs consuming 5W), and throughput capabilities (single-threaded CPU vs. massively parallel GPU). Higher PPP indicates better cost-power-performance trade-off for deployment at scale.
3.9 Optimization Problem Formulation with Hardware Constraints
The complete QEIL optimization synthesizes all objectives with hardware-specific constraints derived from our experimental platform:
$\min_{\mathcal{A}}$ $\quad E_{\text{total}}(\mathcal{A})=\sum_{i}(E_{\text{prefill},i}+E_{\text{decode},i})$ (33)
s.t. $\quad\sum_{l\in L_{i}}M_{l}\leq M_{i}^{\max}$
@@ 3.9 Optimization Problem Formulation with Hardware Constraints
$\quad\tau_{\text{total}}(\mathcal{A})\leq\tau_{\max}$
$\quad C(S,N,T)\geq C_{\min}$
These constraints ensure feasibility on our experimental edge platform: Intel Core Ultra 9 285HX processor (8 cores, 2.80 GHz), 128 GB system RAM (127 GB usable), Intel AI Boost NPU with 20 GB dedicated storage, NVIDIA RTX PRO 5000 Blackwell GPU (96.2 GB total VRAM), and Intel Graphics GPU (72.7 GB shared memory). Memory and power limits reflect realistic edge deployment scenarios where devices operate under strict resource budgets while maintaining inference quality and throughput.
3.10 Experimental Results and Performance Validation
We validate QEIL through comprehensive experiments across five diverse language model families on our heterogeneous edge platform using WikiText-103 benchmark. The evaluation spans models from 125M to 2.6B parameters—GPT-2 (125M), Granite-350M, Qwen2-0.5B, Llama-3.2-1B, and LFM2-2.6B—demonstrating universality across distinct architectures and scales. Our experimental methodology enables intelligent distribution of inference workloads across the platform’s heterogeneous device ecosystem: CPU for embedding and language model head operations (memory-bound, low-compute operations), GPU for compute-intensive prefill phase (high arithmetic intensity), and NPU for memory-bound decode phase (bandwidth-critical autoregressive generation). Each model is evaluated on both standard throughput-optimized and energy-aware orchestration paradigms, enabling direct comparison of performance-energy trade-offs.

### v2

@@ 4.5 Cross-Model Consistency Validation
GPT-2 (125M) +10.5 -47.7 +382
Granite-350M +9.0 -78.2 +460
Qwen2-0.5B +10.5 -46.7 +229
Llama-3.2-1B +7.0 -35.6 +108
LFM2-2.6B +8.0 -35.9 +130
Mean +9.0 -48.8 +262
Std Dev 1.4 17.2 149
LFM2-2.6B IPW computed as energy reduction gain due to coverage-normalized comparison
Conclusion: All ablation findings exhibit consistent patterns across model families, with mean improvements of +9.0pp coverage, -48.8% energy, and substantial IPW gains. The moderate variance in energy reduction (std = 17.2%) reflects model-specific characteristics (mid-range models like Granite-350M benefit most from heterogeneous optimization), while coverage gains remain remarkably stable (std = 1.4pp).
4.6 Energy Consumption and Latency Breakdown Analysis
To provide deeper insight into QEIL’s efficiency gains, we analyze the breakdown of energy consumption and latency across execution modes and hardware configurations.
4.6.1 Total Energy Consumption Comparison
Figure 2 presents the total energy consumption comparison between standard (homogeneous GPU) and energy-aware (heterogeneous QEIL) execution modes across 20 inference samples on GPT-2 (125M).
Figure 2: Total Energy Consumption Comparison between Standard (homogeneous GPU) and Energy-Aware (heterogeneous QEIL) execution modes on GPT-2 (125M) with $S=20$ samples. Standard execution consumes 43,057.7 J while Energy-Aware execution achieves 22,487.8 J, representing a 47.8% reduction in total energy consumption through intelligent heterogeneous orchestration.
Table 7: Detailed Energy Breakdown: Standard vs. Energy-Aware Execution on GPT-2 (125M).
Metric Standard Energy-Aware $\Delta$
Total Energy (J) 43,057.7 22,487.8 -47.8%
Prefill Energy (J) 12,450.2 8,234.1 -33.9%
Decode Energy (J) 28,892.5 12,876.4 -55.4%
Overhead Energy (J) 1,715.0 1,377.3 -19.7%
Avg Power (W) 402.5 83.5 -79.2%
Energy per Token (J) 21.53 11.24 -47.8%
Key Insight: The energy-aware configuration achieves the largest savings during the decode phase (-55.4%), validating our hypothesis that routing memory-bound decode operations to energy-efficient NPUs yields substantial benefits. The prefill phase shows moderate savings (-33.9%) as GPU compute efficiency remains valuable for compute-bound operations.
4.6.2 Latency Breakdown Analysis
Figure 3 presents the latency breakdown comparing CPU-only execution versus heterogeneous CPU-GPU-NPU orchestration.
Figure 3: Latency Breakdown Comparison between CPU-Only and CPU-GPU-NPU (Heterogeneous) execution modes. CPU-Only execution requires 20.7ms total (dominated by compute time at $\sim$18ms), while heterogeneous orchestration achieves 8.6ms total through parallel execution across specialized hardware, representing a 58.5% latency reduction.
Table 8: Detailed Latency Breakdown: CPU-Only vs. Heterogeneous (CPU-GPU-NPU) Execution.
Component CPU-Only (ms) Heterogeneous (ms) $\Delta$
Compute Time 18.2 7.2 -60.4%
Memory Transfer 2.1 0.9 -57.1%
Controller Overhead 0.4 0.5 +25.0%
Total Latency 20.7 8.6 -58.5%
Key Insight: Despite introducing controller overhead (+25% for orchestration coordination), heterogeneous execution achieves 58.5% total latency reduction by leveraging specialized hardware: GPUs for compute-intensive prefill (reducing compute time by 60.4%) and NPUs for memory-efficient decode operations. The memory transfer overhead is also reduced (-57.1%) through optimized data placement across the device hierarchy.
4.6.3 Real-Time Dynamic Orchestrator Visualization
Figure 4 presents a real-time Task Manager snapshot captured during QEIL’s dynamic orchestration, demonstrating the simultaneous utilization of CPU, GPU, and NPU resources during heterogeneous inference execution.
Figure 4: Real-Time Task Manager Visualization of QEIL Dynamic Orchestrator during heterogeneous inference execution on GPT-2 (125M). The snapshot demonstrates simultaneous utilization across multiple processing units: CPU at 9% (2.04 GHz) handling orchestration and lightweight operations, Intel AI Boost NPU at 44% executing memory-bound decode phases, Intel Graphics GPU at 95% processing compute-intensive prefill stages, and NVIDIA RTX PRO 5000 GPU at 21% (57°C) handling overflow compute tasks. This multi-vendor, multi-device parallel execution exemplifies QEIL’s agentic orchestration capabilities in resource-constrained edge environments. Note the GPU temperature of 57°C, well below the 85°C thermal throttling threshold, demonstrating safe thermal operation.
Table 9: Real-Time Device Utilization During QEIL Dynamic Orchestration.
Device Vendor Util. Role
CPU Intel 9% Orchestration, I/O
NPU 0 Intel (AI Boost) 44% Decode (mem-bound)
GPU 0 Intel (Graphics) 95% Prefill (compute-bound)
GPU 1 NVIDIA (RTX 5000) 21% Overflow compute
Memory 20% KV Cache, Activations
Key Insight: The Task Manager snapshot demonstrates QEIL’s ability to orchestrate workloads across multi-vendor hardware simultaneously. Intel GPU handles the majority of compute-intensive prefill at 95% utilization, while Intel NPU processes memory-bound decode at 44%. The NVIDIA GPU serves as overflow capacity at 21%, and the CPU maintains low utilization (9%) while coordinating the entire pipeline. This visualization empirically validates our agentic orchestration framework’s ability to leverage heterogeneous hardware from both the same vendor (Intel CPU + Intel NPU + Intel GPU) and different vendors (Intel + NVIDIA) in a unified, resource-constrained edge deployment.
4.7 Safety and Reliability Validation
To validate QEIL’s safety-first design, we conduct stress tests evaluating thermal protection, fault tolerance, and adversarial robustness.
4.7.1 Thermal Protection Effectiveness
We evaluate thermal protection by running sustained inference workloads and comparing behavior with and without thermal constraints:
Table 10: Thermal Protection Evaluation: 30-minute sustained inference on GPT-2 (125M).
Metric Without Protection With Protection
Max GPU Temp (°C) 89 (throttled) 72
Thermal Throttling Events 47 0
Avg Latency (ms) 1.89 $\pm$ 0.84 1.41 $\pm$ 0.08
Latency 99th Pctl (ms) 4.21 1.58
Total Throughput (tokens) 142,847 156,892
Key Finding: Without thermal protection, the GPU reaches throttling temperatures, causing 47 throttling events with highly variable latency (std = 0.84ms) and reduced total throughput. With thermal protection, the system maintains stable temperatures (72°C peak, well below the 85°C threshold), eliminates throttling events, and achieves higher throughput due to consistent performance. This demonstrates that safety-first design improves rather than degrades overall system performance.
4.7.2 Fault Tolerance and Recovery
We simulate device failures during inference to evaluate fault recovery:
Table 11: Fault Tolerance Evaluation: Recovery from simulated device failures.
Failure Scenario Recovery (ms) Throughput Queries Lost
NPU failure (44% load) 78 -31% 0
GPU failure (95% load) 124 -58% 0
Both GPU failure 156 -72% 0
NPU + 1 GPU failure 98 -64% 0
Key Finding: QEIL achieves zero query loss across all tested failure scenarios, with recovery times under 200ms. Even catastrophic failures (both GPUs failing) result in graceful degradation to CPU-only execution rather than system crash. The throughput impact is proportional to lost compute capacity, demonstrating predictable degradation.
4.7.3 Adversarial Robustness
We evaluate input validation against common attack vectors:
Table 12: Adversarial Robustness: Input validation effectiveness.
Attack Type Blocked System Impact
Oversized input (10× context) 100% None
Malformed UTF-8 100% None
Rapid-fire requests (DDoS) 99.2% 0.8% degradation
Repetition-inducing prompts 94% 6% excess tokens
Key Finding: Input validation successfully blocks malformed and oversized inputs with 100% effectiveness. Rate limiting prevents resource exhaustion from rapid requests. Output sanity checking catches 94% of prompts designed to induce pathological repetition, limiting excess token generation to 6%. These defenses protect against common attack vectors while maintaining normal operation for legitimate requests.
4.8 Cross-Dataset Robustness Evaluation
To demonstrate that QEIL’s benefits generalize beyond WikiText-103, we evaluate on two additional challenging benchmarks: GSM8K (mathematical reasoning) and ARC-Challenge (scientific reasoning). These datasets test fundamentally different capabilities and provide rigorous validation of our framework’s robustness.
4.8.1 GSM8K Mathematical Reasoning Results
GSM8K Cobbe et al. (2021) contains 8,500 grade-school math word problems requiring multi-step reasoning. This benchmark specifically tests whether QEIL’s scaling formalisms extend to complex reasoning tasks where chain-of-thought generation is critical.
Table 13: Cross-Dataset Evaluation on GSM8K (Mathematical Reasoning): QEIL Framework Across Five Model Families. Results demonstrate consistent improvements in energy efficiency and coverage on reasoning-intensive tasks requiring multi-step chain-of-thought generation. All metrics measured with $S=20$ samples and chain-of-thought prompting.
Model Exec Type Accuracy Pass@k Energy IPW Latency PPP
(%) (%) (kJ) (tasks/W) (ms) Score
GPT-2 (125M) Standard 12.4 18.2 52.3 0.045 2.14 8.42
Energy-Aware 15.8 24.6 28.1 0.182 1.67 12.85
Improvement +3.4pp +6.4pp -46.3% +304% -22.0% +52.6%
Granite-350M Standard 18.7 26.4 485.2 0.039 2.08 6.94
Energy-Aware 24.2 35.8 112.6 0.215 1.72 11.26
Improvement +5.5pp +9.4pp -76.8% +451% -17.3% +62.2%
Qwen2-0.5B Standard 24.5 34.2 421.8 0.081 2.18 7.52
Energy-Aware 31.2 44.8 218.4 0.251 1.94 11.89
Improvement +6.7pp +10.6pp -48.2% +210% -11.0% +58.1%
Llama-3.2-1B Standard 35.8 48.6 398.4 0.122 2.34 8.18
Energy-Aware 42.4 58.2 254.8 0.286 1.98 12.45
Improvement +6.6pp +9.6pp -36.0% +134% -15.4% +52.2%
LFM2-2.6B Standard 42.1 56.8 586.2 0.097 2.28 9.84
Energy-Aware 49.8 66.4 372.4 0.178 1.86 14.52
Improvement +7.7pp +9.6pp -36.5% +83% -18.4% +47.6%
Mean Aggregate +6.0pp +9.1pp -48.8% +236% -16.8% +54.5%
GSM8K Key Findings:
  • Consistent coverage improvements: QEIL achieves +9.1pp average pass@k improvement across all models, validating that inference-time scaling extends to mathematical reasoning.
  • Energy efficiency gains preserved: The 48.8% mean energy reduction closely matches WikiText-103 results (48.8%), confirming that our scaling formalisms generalize across task types.
  • Reasoning benefits from heterogeneous orchestration: Larger models (Llama-3.2-1B, LFM2-2.6B) show strong absolute accuracy gains (+6.6-7.7pp), suggesting that chain-of-thought generation benefits from GPU-accelerated prefill and NPU-efficient decode.
4.8.2 ARC-Challenge Scientific Reasoning Results
ARC-Challenge Clark et al. (2018) contains 2,590 difficult science questions requiring commonsense and scientific reasoning. This benchmark tests QEIL’s performance on knowledge-intensive tasks with shorter output sequences.
Table 14: Cross-Dataset Evaluation on ARC-Challenge (Scientific Reasoning): QEIL Framework Across Five Model Families. Results demonstrate robust improvements on knowledge-intensive reasoning tasks with shorter output sequences. All metrics measured with $S=20$ samples.
Model Exec Type Accuracy Pass@k Energy IPW Latency PPP
(%) (%) (kJ) (tasks/W) (ms) Score
GPT-2 (125M) Standard 25.8 34.2 38.6 0.089 1.58 11.24
Energy-Aware 31.4 42.8 19.8 0.398 1.21 16.85
Improvement +5.6pp +8.6pp -48.7% +347% -23.4% +49.9%
Granite-350M Standard 32.4 44.6 358.4 0.090 1.52 9.86
Energy-Aware 39.8 54.2 78.2 0.509 1.28 15.42
Improvement +7.4pp +9.6pp -78.2% +466% -15.8% +56.4%
Qwen2-0.5B Standard 38.2 52.4 312.6 0.122 1.62 10.28
Energy-Aware 46.8 62.8 164.2 0.421 1.48 15.86
Improvement +8.6pp +10.4pp -47.5% +245% -8.6% +54.3%
Llama-3.2-1B Standard 48.6 64.2 294.8 0.165 1.74 11.42
Energy-Aware 56.2 72.8 186.4 0.389 1.52 16.28
Improvement +7.6pp +8.6pp -36.8% +136% -12.6% +42.6%
LFM2-2.6B Standard 54.2 70.4 452.6 0.120 1.68 12.86
Energy-Aware 62.4 78.6 284.8 0.219 1.38 18.24
Improvement +8.2pp +8.2pp -37.1% +83% -17.9% +41.8%
Mean Aggregate +7.5pp +9.1pp -49.7% +255% -15.7% +49.0%
ARC-Challenge Key Findings:
  • Strong accuracy improvements: QEIL achieves +7.5pp mean accuracy improvement and +9.1pp pass@k improvement, demonstrating effectiveness on knowledge-intensive tasks.
  • Consistent energy savings: The 49.7% mean energy reduction aligns closely with WikiText-103 (48.8%) and GSM8K (48.8%), confirming scaling formalism robustness across diverse task types.
  • Latency benefits on shorter sequences: ARC-Challenge uses shorter output sequences than GSM8K, yet QEIL still achieves 15.7% latency reduction, validating that heterogeneous orchestration benefits both long-form reasoning and short-form QA tasks.
4.8.3 Cross-Dataset Consistency Summary
Table 15: Cross-Dataset Consistency: Mean Improvements Across Three Benchmarks.
Metric WikiText GSM8K ARC-C
@@ 4.8.3 Cross-Dataset Consistency Summary
[…]
Std Dev (pp/%) 0.1/0.0 0.1/0.5
[…]
5 Results
We validate QEIL through comprehensive experiments across five diverse language model families on our heterogeneous edge platform using WikiText-103 benchmark. The evaluation spans models from 125M to 2.6B parameters—GPT-2 (125M), Granite-350M, Qwen2-0.5B, Llama-3.2-1B, and LFM2-2.6B—demonstrating generalizability across distinct architectures and scales within the transformer family.
5.1 Inference-Time Scaling Behavior Across Heterogeneous Hardware
We begin by characterizing how inference-time performance scales with sample budget, model size, and hardware configuration. Across five transformer-based model families (125M–2.6B parameters), we observe that inference coverage exhibits consistent power-law scaling with respect to the number of samples. Notably, the inferred scaling exponent remains stable across models and tasks, suggesting that inference-time behavior follows predictable structural patterns analogous to training-time scaling relationships.